Bond Duration: What is the duration of a bond with 7% discount rate, 3-year maturity, and 10% coupon paid annually?
Answer #1
                  K = N
Bond Price =∑ [( Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =3
Bond Price =∑ [(10*1000/100)/(1 + 7/100)^k]     +   1000/(1 + 7/100)^3
                   k=1
Bond Price = 1078.73

phpB5U5Ps.png

Period Cash Flow Discounting factor PV Cash Flow Duration Calc
0 ($1,078.73) =(1+YTM/number of coupon payments in the year)^period =cashflow/discounting factor =PV cashflow*period
1              100.00                                                             1.07                    93.46                  93.46
2              100.00                                                             1.14                    87.34                174.69
3           1,100.00                                                             1.23                  897.93              2,693.78
      Total              2,961.93
Macaulay duration =(∑ Duration calc)/(bond price*number of coupon per year)
=2961.93/(1078.73*1)
=2.745755
Modified duration = Macaulay duration/(1+YTM)
=2.75/(1+0.07)
=2.566127